The edge length of a cubic unit cell is 5.63×10⁻¹⁰m. To find the distance between (111)planes.
1
Expert's answer
2020-12-21T03:53:51-0500

"d_{hkl}=\frac{a}{\sqrt{h^2+k^2+l^2}}"

"d_{111}=\frac{5.63*10^{-10}}{\sqrt{1^2+1^2+1^2}}=3.25*10^{-10}"